# Sliding Speed Influences Electrovibration-Induced Finger Friction Dynamics on Touchscreens

### v1

@@ 3.1 Dynamics of electrovibration-induced finger friction on touchscreens
The estimated gains and cutoff frequencies of the first-order models fitted to the measured friction signal and input message signal, $V_{m}$, under the tested exploration conditions are shown in Figures 6(a) and 6(b). We performed statistical analyses to examine how sliding speed and force influence electrovibration-induced finger friction. In our analysis, the applied force, $F_{n}$, and sliding speed, $\nu$, were treated as independent variables (main effects), while the model parameters – the gain $K$ and cutoff frequency $\omega_{o}$ – were the dependent variables. We assessed the normality of all data using the Kolmogorov-Smirnov test, which indicated that both the gains and cutoff frequencies were not normally distributed ($p<0.01$). Hence, we applied a generalized linear mixed model (GLMM) to test the influence of the independent variables. To account for the variability in these parameters introduced by different participants, the participant ID number was included as a random effect. The resulting mixed model with a linear approximation for the gain parameter $K$ is given as:
@@
Here, $\alpha_{o}$ is the model’s intercept parameter, $\alpha_{1}$, $\alpha_{2}$, and $\alpha_{3}$ are the fixed coefficients corresponding to the effects of force, speed, and the force-speed interaction, $\rho$ is the coefficient for the random between-participant effect, and $\epsilon$ is the model’s residual. The fixed coefficients show the estimated influence of the independent variable on the dependent variable, with other independent variables being constant. Our analysis revealed that the main effects of applied force ($F(1,178)=0.310$, $p=0.578$) and sliding speed ($F(1,214)=0.715$, $p=0.399$), and their interaction ($F(1,210)=0.892$, $p=0.346$) all did not affect the gain. Moreover, between-participant variation also had no significant effect on the measured $K$ values (Estimate: $2.96\times 10^{-5}$, SE: $1.53\times 10^{-5}$, $Z=1.92$, $p=0.054$).
@@
Here, $\beta_{o}$ is the model intercept, $\beta_{1}$, $\beta_{2}$, and $\beta_{3}$ are the fixed coefficients for force, speed, and the force-speed interaction, $\rho$ is the coefficient for the random effect due to participant ID, and $\epsilon$ is the model’s residual. Unlike the statistical analysis results for $K$, the cutoff frequency was significantly affected by the sliding speed ($F(1,211)=25.068$, $p<0.001$). However, also no significant effects of the applied force ($F(1,176)=1.12$, $p=0.291$), the speed-force interaction ($F(1,207)=0.715$, $p=0.399$), and the between-participant effect (Estimate: $4.41\times 10^{4}$, SE: $2.43\times 10^{4}$, $Z=1.816$, $p=0.069$) were found. Hence, only the finger speed significantly affects the cutoff frequency $\omega_{o}$; the corresponding fixed coefficient, $\beta_{2}$, has a value of 13.811 Hz/mm/s, meaning that for every increase in speed by 1 mm/s, the cutoff frequency increases by 13.811 Hz (with respect to an intercept, $\beta_{o}$, of 385.68 Hz). Next to the GLMM analysis, we also performed a post-hoc Wilcoxon signed-rank test to evaluate if the cutoff frequencies across the participants statistically differ with sliding speeds for different forces; see bars at the top of Figure  6(b) for the significantly different pairs. The post-hoc analysis showed a significant difference in cutoff frequency across sliding speeds, particularly between the lower and higher speed ranges. However, as the applied force increased, the cutoff frequency remained insignificant at higher force levels (60–100 mm/s).
$H_{f-e}(s,\nu)=\frac{K\omega_{o}}{s+\omega_{o}(\nu)}\in K=0.0123;\omega_{o}=385.68+13.811\nu$ (7)
Here, $K=0.0123$ is the average gain across all sliding speeds, applied forces, and participants measured in the experiment, as our GLMM analysis showed no effects of these factors on this gain. However, as the cutoff frequency was influenced by the sliding speed, we simplified the corresponding linear model as $\omega_{o}(\nu)=385.68+13.811\nu$ to describe the variation in finger-display interaction with sliding speed. Note that this model is only valid for an input voltage amplitude of 150 $V_{pp}$; it does not capture variations with amplitude changes.
@@ 3.2 Dynamics of finger’s mechanical behavior
Figure 8: Scatter plot showing the Pearson’s correlation between the finger-display interaction parameters—$K$ and $\omega_{o}$—and finger mechanical response parameters—$m$, $b$, $k$. (a) Gain, $K$ vs. finger moving mass, $m$, (b) cutoff frequency, $\omega_{o}$, vs. finger moving mass, $m$, (c) gain, $K$, vs. finger damping, $b$, (d) cutoff frequency, $\omega_{o}$, vs. finger damping, $b$, (e) gain, $K$, vs. finger stiffness, $k$, (f) cutoff frequency, $\omega_{o}$, vs. finger stiffness, $k$. The blue dots indicate the data points, and the dashed line indicates Pearson’s correlation fit. The correlation coefficient, $r$, and significance value, $p$, are shown in the plot.
@@
The mass, damping, and stiffness of the second-order model fitted to the measured finger vibration velocity for input finger friction under different exploration conditions are shown in Figures 7(a)7(b), and 7(c), respectively. Similarly, a Kolmogorov-Smirnov test indicated that these parameters were not normally distributed ($p<0.01$). We assigned the model parameters -— the finger moving mass $m$, finger damping $b$, and finger stiffness $k$ – as dependent variables, the sliding speed and applied force as the main effects, as well as their interaction, and participant ID as a random effect. The GLMM analysis revealed that the finger moving mass was affected by sliding speed ($F(1,205)=7.193$, $p=0.008$) with a fixed coefficient of $3.5\times 10^{-5}$ Kg/mm/s, but not by the applied force ($F(1,176)=0.125$, $p=0.724$) or their interaction ($F(1,201)=1.216$, $p=0.271$). The estimated coefficient values indicate that the finger moving mass increased by $0.0028$ kg from the lowest to the highest sliding speed. The finger damping was not affected by the applied force ($F(1,179)=3.280$, $p=0.071$), sliding speed ($F(1,224)=2.843$, $p=0.093$), or their interaction ($F(1,221)=1.604$, $p=0.207$). The finger stiffness was affected by sliding speed ($F(1,208)=14.465$, $p<0.001$) with a fixed coefficient of $4.036$ N/m/mm/s, but not by the applied force ($F(1,175)=1.148$, $p=0.285$) or their interaction ($F(1,204)=3.79$, $p=0.053$). Based on the obtained coefficient values, the finger stiffness increased by $322.88$ N/m from the lowest to the highest applied force. Moreover, the obtained moving mass (Estimate: $4.99\times 10^{-6}$, SE: $2.41\times 10^{-6}$, $Z=2.068$, $p=0.039$), damping (Estimate: 0.356, SE: 0.180, $Z=1.979$, $p=0.048$), and stiffness (Estimate: $2.36\times 10^{4}$, SE: $1.16\times 10^{4}$, $Z=2.03$, $p=0.042$) values were significantly affected by the participant ID. We also conducted a post-hoc Wilcoxon signed-rank test to reveal for which tested sliding speeds these values significantly differ (see bars above figures in Figures 7(a)7(b), and 7(c)).
@@ 3.3 Influence of skin mechanics on finger-display interaction
As a final analysis step, we performed a Pearson’s correlation analysis between the finger-display interaction parameters (the gain, $K$, and cutoff frequency, $\omega_{o}$) and the finger mechanical response parameters (the moving mass, $m$, damping, $b$, and stiffness, $k$) for all the sliding speeds and force, as shown in Figure 8. The results indicate that gain significantly decreased with an increase in finger moving mass, and increased with an increase in finger damping. Similarly, the cutoff frequency significantly increased with higher finger moving mass and finger stiffness.
@@ 4 Discussion
In this study, we investigated how exploration conditions influence electrovibration-induced finger friction on touchscreens and the role of skin mechanics in this relation. We measured the contact forces and skin accelerations generated as ten participants moved their index fingers on an electrovibration display under varying exploratory forces and speeds. Using this data, we developed mathematical models to characterize the electrovibration-induced finger friction and the finger vibrations it causes. Finally, we analyzed how the models’ parameters change across different sliding speeds and applied force settings.
Our results demonstrate that the cutoff frequency of electrovibration-induced friction response increased with sliding speed (Figure 6(b)). We attribute this increase to changes in skin mechanics with sliding speed. In our measurements, both the finger moving mass and stiffness increased with sliding speed and consequently influenced the natural frequency of the finger dynamics, $\omega_{finger}=\sqrt{{k}/{m}}$; check Figure 7(a) and 7(c). While higher stiffness raises the resonance frequency, the increased moving mass lowers it, potentially reducing the cutoff frequency. However, our GLMM analysis revealed that the variation in the stiffness was stronger than the change in mass, resulting in an overall increase in the cutoff frequency. For every 1 mm/s increase in sliding speed, the predicted finger stiffness increased by 4.036 N/m while the finger moving mass increased by $3.223\times 10^{-5}$ kg. Dividing these changes by their measured values at lower speeds, they correspond to increases of 0.82% in mass and 1.11% in stiffness. Additionally, Figure 8) indicates that the correlation coefficient of finger stiffness with cutoff frequency is 30% higher than the correlation with the finger moving mass. These observations provide further evidence that the finger acts less elastic with sliding speed, causing an increase in the cutoff frequency.
@@
Unlike the cutoff frequency, our results (Figure 6(a)) showed that the gain of the electrovibration-induced finger friction was not affected by the sliding speed. We explain this invariance with the observed simultaneous variation in finger moving mass and damping. Our correlation results revealed that the gain decreased with increasing finger moving mass, but increased with increasing finger damping (see Figure 8(a) and Figure 8(c)). A higher finger moving mass implies a lower finger vibration amplitude at low frequencies, which is consistent with our results. Although the finger moving mass increased with speed, the increase is $3.223\times 10^{-5}$ kg per mm/s speed, which is small enough that it does not drastically affect the gain. Similar to the finger moving mass, an increase in finger damping reduces the finger vibration intensity at the resonance. However, our results showed that, like the gain, finger viscosity also remained constant across the sliding speed.
@@
Our results showed that applied force did not have a significant effect on the electrovibration-induced friction and on the finger vibration caused by it (see Figures 6(a)6(b)7(a)7(b), and  7(c)). Interestingly, previous studies have reported an increase in generated friction force due to electrovibration with applied force [27]. Similarly, earlier studies [15, 44, 16] observed a rise in the finger’s mechanical parameters, such as mass, damping, and stiffness with applied force. This contradiction between our results and earlier studies could be caused by the low force ranges for the active applied forces ($0.2\sim 0.6$ N) in our study  [45]. Vardar and Kuchenbecker [13] also found no significant change in the relationship between the input voltage and electrovibration force with varying applied force. They argued that normal forces at 0.5 and 1.0 N still represent sufficiently light and moderate touches not to cause any change in the skin mechanics, which is consistent with our results.
Our findings revealed significant variability in finger skin mechanics among participants, likely due to differences in finger and skin biomechanics, skin moisture, and finger size [46, 47, 48, 49, 50]. Interestingly, the finger electrovibration response parameters ($K$ and $\omega_{o}$) did not show a significant between-participant effect. Since the finger’s mechanical response reflects dynamics across the entire fingertip (from fingerpad to fingernail), while the friction response focuses only on contact changes, it might be concluded that the contact variations were not substantial enough to be detected in the analysis. Also, the skin mechanics and finger-electrovibration friction models used in this study likely oversimplify the complex finger-electrovibration interaction, meaning random effects may not influence them equally.
@@
Our results indicate that at low applied forces ($0.2$ N to $0.6$ N), the spectrum for the electrovibration-induced finger friction remains unchanged. However, as sliding speed increases, the higher-frequency signals of the friction spectrum experience less attenuation, making these frequencies more pronounced. If this speed-dependent behavior is not accounted for in electrovibration-based texture rendering, users may perceive virtual textures differently from the intended ones (e.g., finger friction between real textures or a designed virtual slider in a car cockpit). To ensure a more realistic and robust experience, texture rendering algorithms should dynamically adjust any high-frequency signal based on the sliding speed. With these insights, we proposed an empirically derived model (Equation (7)) that can be used to predict the evolution of finger friction generated by electrovibration based on finger sliding speed for the given input signal magnitude. Using this knowledge, researchers can incorporate the corresponding inverse electrovibration-induced finger friction filter in texture rendering algorithms that continuously minimize the error between the intended texture spectrum and the generated friction spectrum, dynamically adjusting for the input voltage to output friction variations for the finger electrovibration exploration behavior. In the future, we aim to use the proposed finger friction filter and develop a closed-loop texture rendering algorithm that adapts to user exploration, enabling a more natural and free texture exploration experience. Moreover, we aim to evaluate the practical electromechanical response to test whether participants can perceptually differentiate it from user- and condition-specific compensations.
Despite our careful investigation, our work still has limitations. We asked participants to perform a constrained active finger exploration by following the cues in the experimental GUI. This task was quite difficult for some participants, and we had to remove recorded data from seven participants. The participants also found maintaining a consistent contact angle and contact area challenging. Also, the participants performed the experiments in multiple sessions, which could have added to the variability in the model parameters. The potential variations in the contact parameters may have influenced the air gap between the finger and the screen and, consequently, the measured forces. This variability could also explain the lack of a significant trend between gain and sliding speed. It can be reduced when the participant’s finger exploration is controlled with motorized linear stages in the experimental setup.

### v2

@@ 3.1 Dynamics of electrovibration-induced finger friction on touchscreens
Figure 6: The obtained transfer functions, $H_{f-e}$, representing the relationship between the measured friction force, $F_{f}$, and the input message signal, $V_{m}$, across tested exploration speeds and forces for all participants. The color-coded center lines show the mean transfer function across different speeds, while the shaded region represents the standard error across participants. The dashed vertical lines mark the 3-dB cutoff frequencies, $\omega_{o}$.
@@
The estimated first-order transfer functions, which represent the relationship between the measured friction force, $F_{f}$, and the input message signal, $V_{m}$, across different applied forces and sliding speeds, are shown in Figure 6. The corresponding estimated model gains and cutoff frequencies are shown in Figures 7(a) and 7(b).
We performed statistical analyses to examine how sliding speed and force influence electrovibration-induced finger friction. In our analysis, the applied force, $F_{n}$, and sliding speed, $\nu$, were treated as independent variables (main effects), while the model parameters – the gain $K$ and cutoff frequency $\omega_{o}$ – were the dependent variables. We assessed the normality of all data using the Kolmogorov-Smirnov test, which indicated that both the gains and cutoff frequencies were not normally distributed ($p<0.01$). Hence, we applied a generalized linear mixed model (GLMM) to test the influence of the independent variables. To account for the variability in these parameters introduced by different participants, the participant ID number was included as a random effect. The resulting mixed model with a linear approximation for the gain parameter $K$ is given as:
@@
Here, $\alpha_{o}$ is the model’s intercept parameter, $\alpha_{1}$, $\alpha_{2}$, and $\alpha_{3}$ are the fixed coefficients corresponding to the effects of force, speed, and the force-speed interaction, $\rho$ is the coefficient for the random between-participant effect, and $\epsilon$ is the model’s residual. The fixed coefficients show the estimated influence of the independent variable on the dependent variable, with other independent variables being constant. Our analysis revealed that the main effects of applied force ($F(1,178)=0.310$, $p=0.578$) and sliding speed ($F(1,214)=0.715$, $p=0.399$), and their interaction ($F(1,210)=0.892$, $p=0.346$) all did not affect the gain.
To evaluate the variability across participants, we conducted a likelihood ratio test (LRTLRT[43] by comparing the full GLMM model, which included the random effect, with a reduced model excluding it. This procedure involved calculating the -2 log-likelihood values for both the full ($2LL_{full}$) and reduced ($2LL_{reduced}$) models, defined as:
$LRT=-2LL_{reduced}-(-2LL_{full})$ (6)
If the resulting $LRT$ value exceeds the critical value from the chi-square distribution (with degrees of freedom equal to the number of random effects, here $df=1$), the random effect is considered statistically significant. In our analysis, $2LL_{reduced}=-1594.327$ and $2LL_{full}=-1631.082$, yielding an $LRT$ value of 36.755. This number exceeds the critical chi-square value of 3.84 at $p=0.05$, indicating that the random effect due to participants had a statistically significant impact on the model.
@@
Here, $\beta_{o}$ is the model intercept, $\beta_{1}$, $\beta_{2}$, and $\beta_{3}$ are the fixed coefficients for force, speed, and the force-speed interaction, $\rho$ is the coefficient for the random effect due to participant ID, and $\epsilon$ is the model’s residual. Unlike the statistical analysis results for $K$, the cutoff frequency was significantly affected by the sliding speed ($F(1,211)=25.068$, $p<0.001$). Moreover, the random effect due to participants was also statistically significant ($2LL_{reduced}=3732.094$ and $2LL_{full}=3698.635$, yielding an $LRT$ = 33.459 >> 3.84). However, no significant effects of the applied force ($F(1,176)=1.12$, $p=0.291$), and the speed-force interaction ($F(1,207)=0.715$, $p=0.399$) were found.
These analyses show that only the finger speed significantly affects the cutoff frequency $\omega_{o}$ with a corresponding fixed coefficient, $\beta_{2}$, value of 13.811 Hz/mm/s. This result indicates that for every 1 mm/s increase in speed, the cutoff frequency increases by 13.811 Hz, relative to an intercept, $\beta_{o}$, of 385.68 Hz. Multiplying this coefficient by the difference between the lowest and highest speeds yields a total increase of 1104 Hz. Additionally, dividing the coefficient by the mean cutoff frequency at the lowest speed indicates a 2.47% increase in cutoff frequency for each unit increase in speed.
Next to the GLMM analysis, we also performed a post-hoc Wilcoxon signed-rank test to evaluate if the cutoff frequencies across the participants statistically differ with sliding speeds for different forces; see bars at the top of Figure 7(b) for the significantly different pairs. The post-hoc analysis showed a significant difference in cutoff frequency across sliding speeds, particularly between the lower and higher speed ranges. However, as the applied force increased, the cutoff frequency remained insignificant at higher force levels (60–100 mm/s).
$H_{f-e}(s,\nu)=\frac{K\omega_{o}(\nu)}{s+\omega_{o}(\nu)}\in K=0.0123;\omega_{o}=385.68+13.811\nu$ (8)
Here, $K=0.0123$ represents the average gain across all sliding speeds and applied forces measured in the experiment, as our GLMM analysis showed no significant effects of these factors on the gain. However, as the cutoff frequency was influenced by the sliding speed, we simplified the corresponding model as $\omega_{o}(\nu)=385.68+13.811\nu$, capturing the variation in finger-display interaction with sliding speed. Note that this model is only valid for an input voltage amplitude of 150 $V_{pp}$; it does not capture variations with amplitude changes. Additionally, it reflects an average response, omitting inter-participant variability.
@@ 3.2 Dynamics of finger’s mechanical behavior
The mass, damping, and stiffness of the second-order model—fitted to the measured finger vibration velocity driven by the electrovibration-induced friction input—are presented in Figures 8(a)8(b), and 8(c), corresponding to sliding speed and applied force conditions. A Kolmogorov-Smirnov test confirmed that these parameters were not normally distributed ($p<0.01$). Accordingly, we assigned the model parameters—the finger moving mass $m$, damping $b$, and stiffness kk—as dependent variables, the sliding speed and applied force as the main effects, as well as their interaction, and participant ID as a random effect.
The GLMM analysis revealed that the finger moving mass was affected by sliding speed ($F(1,205)=7.193$, $p=0.008$) with a fixed coefficient of $3.5\times 10^{-5}$ Kg/mm/s, but not by the applied force ($F(1,176)=0.125$, $p=0.724$) or their interaction ($F(1,201)=1.216$, $p=0.271$). The estimated coefficient indicates an increase of approximately $0.0028$ kg in the moving mass from the lowest to the highest sliding speed. Normalizing this value with respect to the moving mass from the lowest speed, we observed a 0.82% increase in moving mass for a unit increase in sliding speed.
The finger damping was not affected by the applied force ($F(1,179)=3.280$, $p=0.071$), sliding speed ($F(1,224)=2.843$, $p=0.093$), or their interaction ($F(1,221)=1.604$, $p=0.207$).
The finger stiffness was affected by sliding speed ($F(1,208)=14.465$, $p<0.001$) with a fixed coefficient of $4.036$ N/m/mm/s, but not by the applied force ($F(1,175)=1.148$, $p=0.285$) or their interaction ($F(1,204)=3.79$, $p=0.053$). Based on the obtained coefficient values, the finger stiffness increased by approximately $322.88$ N/m from the lowest to the highest sliding speed. When normalized relative to the stiffness at the lowest speed, this value corresponds to a 1.11% increase in stiffness per unit increase in sliding speed.
Moreover, the obtained moving mass ($2LL_{reduced}=-2267.563$ and $2LL_{full}=-2352.209$, yielding an $LRT$ = 84.646 >> 3.84), damping ($2LL_{reduced}=592.744$ and $2LL_{full}=550.981$, yielding an $LRT$ = 41.763 >> 3.84), and stiffness ($2LL_{reduced}=3301.393$ and $2LL_{full}=3241.765$, yielding an $LRT$ = 59.628 >> 3.84) values were significantly affected by the participant ID. Additionally, a post-hoc Wilcoxon signed-rank test was performed to determine which sliding speed conditions produced statistically significant differences in these skin mechanics parameters. Significant comparisons are indicated by the bars above the data in Figures 8(a)8(b), and 8(c).
Figure 9: Scatter plot showing the Spearman’s correlation between the cutoff frequency, $\omega_{o}$, of the electrovibration-induced finger friction response and finger stiffness, $k$. Colored dots represent individual data points for each participant, with corresponding dashed lines indicating participant-specific correlation fit. The Spearman correlation coefficient, $r$, and the associated significance value, $p$, are displayed on the plot. The overall correlation across all participants is represented by the solid black line.
3.3 Percentage contribution of applied force and sliding speed
Based on the GLMM coefficients, we computed the percentage contributions of applied force and sliding speed at preferred exploration conditions of $F_{n}=0.4$ N and $\nu=80$ mms/s [37] to the electrovibration-induced finger friction parameters ($K$ and $\omega_{o}$) and skin mechanics parameters ($m$, $b$, and $k$) as summarized in Table 1. For details regarding the calculation procedure, please refer to E.
Table 1: Percentage contribution of finger-electrovibration interaction friction response parameters and skin’s mechanical parameters.
Parameter
Force
coefficient
Speed
coefficient
%Force %Speed
Electrovibration-induced
finger friction
response
Gain, $K$ -0.005 -4.581 $\times 10^{-5}$ 99.1 0.9
Cut off
frequency , $\omega$
1.121 25.068 18.85 87.35
Skin mechanics Mass, $m$ 0.125 7.193 12.64 87.35
Damping , $b$ 3.280 2.843 83.02 16.97
Stiffness , $k$ 1.148 14.465 31.22 68.77
@@ 3.4 Influence of skin mechanics on finger-display interaction
Finally, we performed a Spearman’s correlation [44] analysis between the finger-display interaction parameters (the gain, $K$, and cutoff frequency, $\omega_{o}$) and the finger mechanical response parameters (the moving mass, $m$, damping, $b$, and stiffness, $k$) across all sliding speeds and force, as shown in Figure 15. The results indicate that the gain decreased with increasing finger moving mass, and increased with increasing finger damping. Similarly, the cutoff frequency increased with higher finger moving mass. Among all parameter pairs, only the cutoff frequency and finger stiffness exhibited a low positive correlation ($r=0.34,p=2.8\times 10^{-8}$, see Figure 9). Furthermore, seven out of ten participants showed a correlation greater than 0.52 ($p<0.009$)between their finger stiffness and cutoff frequency.
@@ 4 Discussion
In this study, we investigated how exploration conditions influence electrovibration-induced finger friction on touchscreens and the role of skin mechanics in this relation, with particular focus on texture rendering. We measured the contact forces and skin accelerations generated as ten participants moved their index fingers on an electrovibration display under varying exploratory forces and speeds. Using this data, we developed mathematical models to characterize the electrovibration-induced finger friction and the finger vibrations it causes. Finally, we analyzed how the models’ parameters change across different sliding speeds and applied force settings.
Our results demonstrate that the cutoff frequency of electrovibration-induced finger friction response increased with sliding speed (Figures 6 and 7(b)). We attribute this increase to changes in skin mechanics with sliding speed. In our measurements, both the finger moving mass and stiffness increased with sliding speed and consequently influenced the natural frequency of the finger dynamics, $\omega_{finger}=\sqrt{{k}/{m}}$; check Figure 8(a) and 8(c). While higher stiffness raises the resonance frequency, the increased moving mass lowers it, potentially reducing the cutoff frequency. However, our GLMM analysis revealed that the variation in the stiffness was stronger than the change in mass, resulting in an overall increase in the cutoff frequency. For every 1 mm/s increase in sliding speed, the predicted finger stiffness increased by 1.11% while the finger moving mass increased by 0.82%. Furthermore, we observed statistically significant correlations between estimated cutoff frequency and finger stiffness (Figure 9), reinforcing the role of stiffness as the dominant factor. These findings indicate that with increasing speed, the finger exhibits less elastic behavior, causing an increase in the cutoff frequency.
@@
Unlike the cutoff frequency, our results (Figures 6 and  7(a)) showed that the gain of the electrovibration-induced finger friction was not affected by the sliding speed. We explain this invariance with the observed simultaneous variation in finger moving mass and damping. Our correlation results revealed that the gain decreased with increasing finger moving mass, but increased with increasing finger damping (see Figure 15(a) and Figure 15(c)). A higher finger moving mass implies a lower finger vibration amplitude at low frequencies, which is consistent with our results. Although the finger moving mass increased with speed, the increase is $3.223\times 10^{-5}$ kg per mm/s speed, which is small enough that it does not drastically affect the gain. Similar to the finger moving mass, an increase in finger damping reduces the finger vibration intensity at the resonance. However, our results showed that, like the gain, finger viscosity also remained constant across the sliding speed.
@@
Our results showed that applied force did not have a significant effect on the electrovibration-induced friction and on the finger vibration caused by it (see Figures, 67(a)7(b)8(a)8(b), and  8(c)). Interestingly, previous studies have reported an increase in generated friction force due to electrovibration with applied force [27]. Similarly, earlier studies [15, 52, 16] observed a rise in the finger’s mechanical parameters, such as mass, damping, and stiffness with applied force. This contradiction between our results and earlier studies could be caused by the low force ranges for the active applied forces ($0.2\sim 0.6$ N) in our study  [53]. Vardar and Kuchenbecker [6] also found no significant change in the relationship between the input voltage and electrovibration force with varying applied force. They argued that normal forces at 0.5 and 1.0 N still represent sufficiently light and moderate touches not to cause any change in the skin mechanics, which is consistent with our results.
Our findings revealed substantial inter-participant variability in electrovibration-induced finger friction (Figures 67(a), and 7(b)) and skin mechanics (Figures 8(a), 8(b), and 8(c)). These differences were statistically significant, as confirmed by likelihood ratio tests on the GLMM models, which showed that the electrovibration-induced finger friction response parameters ($K$ and $\omega_{o}$) and skin mechanics parameters ($m$, $b$, and $k$) significantly varied between participants. Similar inter-participant variability has been reported in previous studies on the finger-electrovibration interaction studies  [10, 27, 54]. Such variability is likely driven by individual differences in finger and skin biomechanics, skin hydration, and finger size [55, 56, 16, 46, 57, 58, 59, 60]. These findings suggest that the design of haptic feedback in electrovibration displays may benefit from accounting for individual differences in skin properties.
@@
Our results indicate that at low applied forces ($0.2$ N to $0.6$ N), the spectrum for the electrovibration-induced finger friction remains unchanged. However, as sliding speed increases, the higher-frequency signals of the friction spectrum experience less attenuation, making these frequencies more pronounced. If this speed-dependent behavior is not accounted for in electrovibration-based texture rendering, users may perceive virtual textures differently from the intended ones (e.g., finger friction between real textures or a designed virtual slider in a car cockpit). To ensure a more realistic and robust experience, texture rendering algorithms should dynamically adjust any high-frequency signal based on the sliding speed.
Based on these findings, we proposed an empirically derived model (Equation (8)) that predicts the evolution of finger friction generated by electrovibration based on finger sliding speed for the given input signal magnitude. This model enables researchers to implement a corresponding inverse filter within texture rendering algorithms to minimize the discrepancy between the intended texture spectrum and the actual friction spectrum. The filter dynamically adjusts input voltage to produce the desired frictional response during finger interaction with electrovibration surfaces. In future work, we plan to integrate this filter into a closed-loop texture rendering algorithm that adapts in real time to user exploration, supporting a more natural and freeform tactile interaction experience.
Notably, our model (Equation (8)) is based on parameters averaged across participants, following a methodology similar to that of Felicetti et al. [61], who developed a generalized model by averaging the transfer functions of input voltage to output vibration across users in a vibrotactile device for texture rendering. We plan to evaluate the perceptual consequences of using averaged electrovibration-induced finger friction response versus participant- or condition-specific compensations.
Despite our careful investigation, our work still has limitations. Participants were instructed to perform a constrained active finger exploration by following the cues in the experimental GUI. However, this task proved difficult for some individuals, resulting in the exclusion of data from seven participants due to inconsistencies or errors in their task execution. Participants reported difficulty maintaining a consistent contact angle, applied force, and sliding speed during exploration. Such variations likely influenced the finger-contact behavior, thereby affecting the measured interaction forces [6]. The experiments were also conducted over multiple sessions, which may have introduced further variability in the estimated model parameters. This variability is further reflected in the large standard errors observed in Figure 6, particularly at higher applied forces and lower sliding speeds, indicating difficulties in maintaining consistent exploration behavior. Although the trial rejection method described in C helped ensure data reliability, a more controlled approach, such as passive finger exploration, could improve consistency. In future work, we plan to address this limitation by developing a motorized linear-stage setup to standardize finger exploration and enable data collection from a broader participant population.
Moreover, we employed simplified models to represent electrovibration-induced finger friction (a first-order system model) and skin’s mechanical response (a second-order mass-spring-damper model), treating them separately. While higher-order models that integrate both responses are feasible, we deliberately adopted this simplified approach for texture rendering applications on electrovibration displays. As previously discussed, texture rendering on electrovibration displays relies on modulating friction by applying voltage signals, which must be calibrated to match the exploration condition for robust results, as evidenced by our results. Using a higher-order models to represent the full interaction dynamics could lead to overcompensation during real-time rendering. Additionally, to better understand the effect of skin mechanics on the estimated finger-display interaction parameters, we opted for a one-degree-of-freedom (DOF) second-order model. While a multi-DOF model would more accurately capture the complexity of skin behavior, its implementation would require measurements of vibration velocities at multiple locations on the finger, such as the contact point and the nearby regions, accounting for the motion of independently vibrating mechanical elements. However, such a detailed understanding of internal dynamics may not necessarily provide additional insight into how model parameters influence the finger-display interaction response, nor significantly enhance rendering fidelity.
Nevertheless, we acknowledge that the simplified nature of our models may overlook complex interfacial phenomena, such as additional damping and stiffness effects, which are important for advancing the scientific understanding of finger-surface interactions. As part of future work, we aim to develop a more comprehensive multi-DOF model that more accurately captures these dynamics. For this goal, we will measure multiple independent vibration vectors between the contact point and the finger nail, enabling a 2-DOF or multi-DOF representation. Currently, our skin mechanics model is limited by its reliance on a single vibration signal, measured as acceleration at the fingernail.